OpenStudy (muzzack):

Solve for x: 2 over 5 (x - 2) = 4x. 2 over 9 9 negative 2 over 9 negative 9 over 2

OpenStudy (muzzack):

@Monster1234

OpenStudy (muzzack):

\[\frac{ 2 }{ 5 }(x-2)=4x\]

OpenStudy (anonymous):

multiply both sides by 5 and bring x to one side and constants to other side

OpenStudy (muzzack):

\[\frac{ 2 }{ 5 }=4x+\frac{ 4 }{ 5}\] im guessing

OpenStudy (muzzack):

sorry forgot to add 2/5 x

OpenStudy (anonymous):

2(x-2)=20x

OpenStudy (muzzack):

ohh and what else

OpenStudy (anonymous):

2x-4=20x 18x=-4 solve for x